# DP-2Stage: Adapting Language Models as Differentially Private Tabular Data Generators

### v1

@@ 1 Introduction
In this work, we demonstrate that directly fine-tuning LLMs under DP constraints leads to sub-optimal performance. Our findings, shown in Figure 3, reveal that LLMs struggle to generate coherent and structured text when fine-tuned with DP. This challenge stems from inefficient privacy budget allocation to potentially non-private elements, such as table structures or non-functional tokens.
@@
We evaluated DP-2Stage on two datasets and observed improvements over standard DP fine-tuning across the metrics considered. Notably, DP-2Stage-U exhibited faster inference times (up to 21x) compared to DP-2Stage-O and DP-Standard. Additionally, we observe that fine-tuning with DP while avoiding column shuffling, as used in  Borisov et al. (2023), yielded better performance under DP constraints but underperformed in Non-DP scenarios. Meanwhile, DP benchmarks considered in this work (e.g., DP-CTGAN) demonstrated competitive performance on utility-based metrics but underperformed in marginal-based metrics compared to our proposed DP-2Stage framework.
@@
  • Achieved 3-7% relative reductions in perplexity and 1-2% relative improvements in marginal-based metric (averaged over two datasets) compared to standard DP fine-tuning, with experiments repeated five times per model and four synthetic datasets generated per model. Additionally, DP-2Stage-U demonstrated significantly faster inference—up to 21x faster than DP-Standard and DP-2Stage-O.
@@ 2 Related work
Language models have been extensively studied over the years, evolving from statistical models (Jelinek, 1998) and recurrent neural networks (Hochreiter & Schmidhuber, 1997) to the latest transformer-based architectures (Vaswani et al., 2017). These advancements, supported by the attention mechanisms and rich text datasets, have led to the emergence of large-scale language models (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023; Achiam et al., 2023), a new generation of language models. These models, together with various fine-tuning strategies (Hu et al., 2021; Zhou et al., 2022), have enabled several interesting applications (Borisov et al., 2023; Wang & Fritz, 2024). Borisov et al. (2023) proposed using pre-trained LLMs for synthesizing tabular data in a non-private setting by converting tabular data into text-like formats, significantly improving the performance and paving a new path toward LLM-based tabular generators. In light of the potential, our work delves into applying these advanced models to a relatively unexplored domain: the generation of tabular data under differential privacy constraints. While previous studies (Yu et al., 2021; Li et al., 2022) have explored private fine-tuning of LLM, our focus specifically targets tabular data generation, a unique challenge that diverges from the broader scope of text comprehension. Recently, a concurrent work (Tran & Xiong, 2024) shared a similar spirit of two-stage training. However, this work does not provide insight into the impact of pseudo data used in the first stage and examines a different family of LLM and fine-tuning strategy compared to this work. Our focus is on GPT-2, as utilized in the pioneering work by Borisov et al. (2023), to gain deeper insights into how effective design choices in non-differentially private (Non-DP) settings translate to DP contexts. We offer a detailed analysis of pseudo-data choice, the impact of column shuffling on generation of coherent synthetic tabular data, identify the limitations of the two-stage approach, and discuss open challenges.
@@ 3.1 Language Models for Tabular Data
Let $f$ represent template, which defines the general pattern for organizing tabular data. Using GReaT serialization (Borisov et al., 2023), the template is expressed as $f=``\texttt{<key> is <value>,}^{\prime\prime}$, specifying how keys ($k\in{\mathcal{K}}$), values ($v\in{\mathcal{V}}$), and non-functional elements ($c\in{\mathcal{C}}$, such as “is” and “,”) combine to form a record. An input dataset denoted as ${\mathcal{S}}$, is a realized instance of this $f$, instantiated with tokens $(k,v)$ from a record ${\bm{w}}=({\mathcal{K}},{\mathcal{V}})$. For example, with $k=\{\texttt{age}\}$ and $v=\{32\}$, ${\mathcal{S}}$ is instantiated as $f(k,v)=``\texttt{age is 32,}^{\prime\prime}$.
@@
Column shuffling involves randomly altering the order of column-aligned values within each batch during training. For example, given columns age, sex, and income with values 32, female, and >50k. A detailed illustration is provided in Figure 2. This mechanism happens at every iteration and has been shown to effectively prevent the model from relying on spurious dependency in Non-DP settings (Borisov et al., 2023). However, it complicates DP training due to gradient perturbations, often resulting in higher perplexity and different behavior compared to Non-DP models, as shown in Figure 3. Disabling shuffling fixes the order across iterations.
@@
Figure 2: Illustration of column shuffling. The order of entries is permuted in each iteration.
@@ 3.2 Differential Privacy
Figure 3: DP-2Stage (Ours) vs. Standard DP fine-tuning on the Adult dataset with $\varepsilon=1$. DP-2Stage-O refers to the stage 2 model fine-tuned using out-distribution pseudo data (Airline dataset) in stage 1, while DP-2Stage-U is the stage 2 model fine-tuned using data sampled independently from a Uniform distribution, with statistics derived from the Adult dataset as the pseudo-dataset in stage 1. Perplexity results are displayed from left to right for all words, values, keys, and non-functional words (c.f. section 3.1). The top plot shows model trained without column shuffling, and the bottom shows model trained with column shuffling. Total and Value Perplexity for top and bottom plots have fixed y-axis for ease of comparison.
@@ 4 DP-2Stage
While large language models (LLMs) demonstrate impressive capabilities in generating synthetic tabular data in non-differentially private (Non-DP) settings, as we will later show, fine-tuning them under DP constraints remains challenging for tabular data generation. As illustrated in Figure 3 (bottom), standard DP training, where the LLM model is directly fine-tuned using DPSGD (Abadi et al., 2016), often results in suboptimal performance in terms of the perplexity metric. We hypothesize that this is due to the inefficient allocation of the privacy budget during the learning of data structures. Models employing this approach struggle to reduce the perplexity of keys (e.g., column names) and other non-functional words, despite these elements typically posing minimal privacy risks. Consequently, applying DP to such components not only wastes the privacy budget but also diminishes the overall utility of the model.
@@ 5.1 Implementation Details
We utilized the GPT-2 Radford et al. (2019) model from Hugging Face [^1] https://huggingface.co/ and integrated Opacus [^2] https://opacus.ai/ for differential privacy (DP) training, leveraging the BatchMemoryManager for efficient micro-batching. We fully fine-tuned the model with DP-Adam optimizer using a learning rate of $5e^{-5}$ and a linear scheduler, setting the maximum gradient norm $C=1$ and a target delta $\delta=1e^{-5}$. In the second stage of DP-2Stage, the parameter $\lambda$ was consistently set to 0.65 across all experiments. To convert the table-to-text, we used the GReaT serialization method (Borisov et al., 2023). Unlike GReaT, we maintained a fixed column ordering for each sampled dataset for the DP approaches. The Non-DP benchmark using GPT-2, the DP-Standard model, and the stage 2 models of DP-2Stage are all trained for 10 epochs, while the stage 1 model of DP-2Stage is trained for 5 epochs.
@@
To sample from the model, we condition on the target token for each dataset. For example, in the Adult dataset, where “income” is the target token, sampling begins with the prompt ‘income is ’, and generation continues until all columns have been sampled. A rejection sampling approach was used to discard incomplete generations. However, for the approach with column shuffling enabled (Table 4), rejection sampling proved to be very slow and sometimes failed to complete any generation. In these cases, we employed an imputation-based approach, as detailed in Appendix B.1.
@@
We compare our approach to Non-LLM-based methods under DP and Non-DP settings. For Non-DP baselines, we considered CTGAN, TVAE (Xu et al., 2019) and VAE (adaptation of TVAE using standardized preprocessing technique for the numerical columns i.e removing the mean and scaling to unit variance) while DP baselines include DP-CTGAN, DP-GAN, and DP-VAE. For CTGAN and TVAE, we followed the implementation [^3] https://github.com/sdv-dev/CTGAN, while DP-CTGAN and DP-GAN were implemented using the SmartNoise SDK [^4] https://github.com/opendp/smartnoise-sdk. The numerical column preprocessing was conducted with $\varepsilon=0.1$, while the remaining privacy budget was reserved for private training. For DP-VAE, we adapted the VAE model using SmartNoise preprocessing with $\varepsilon=0.1$.
All experiments were conducted using NVIDIA A100-SXM4-40GB GPUs.
Datasets.
We evaluated two tabular datasets: the Adult Income dataset, which contains over 30,000 samples and 15 attributes used to predict whether an individual’s annual income exceeds \$50,000, hosted on the UCI Machine Learning Repository [^5] https://archive.ics.uci.edu/dataset/2/adult; and the Airline Passenger Satisfaction dataset, which includes over 100,000 samples and 24 attributes related to passenger satisfaction, available on Kaggle [^6] https://www.kaggle.com/datasets/teejmahal20/airline-passenger-satisfaction. Unlike the more commonly benchmarked Adult dataset, the Airline dataset offers a fresh perspective on the ability of deep learning models to capture complex relationships within diverse tabular data and large sample sizes. Table 1 summarizes the statistics of the tabular data. More details can be found in Appendix D.1.
# Train # Test # N # C # Total
Adult 30932 16858 6 9 15
Airline 103904 24976 19 5 24
@@
When training DP-2Stage on the Adult dataset, which is considered as the private dataset to be protected, the publicly available Airline dataset is used as the out-distribution pseudo data for training the Stage 1 model in the DP-2Stage-O approach, and similarly for the reverse scenario.
Evaluation Metrics.
We evaluate the quality of synthetic tabular datasets using two prominent metrics, as detailed in Appendix C, based on the held-out test dataset: Machine Learning Efficacy (utility) and the Normalized Histogram Intersection (fidelity). Machine Learning Efficacy assesses the utility of synthetic data by comparing the performance of machine learning models trained on synthetic data to those trained on real data. The Normalized Histogram Intersection quantifies fidelity by measuring the similarity between the marginal distributions of real and synthetic data, summing the minimum probabilities across corresponding bins, and averaging across all columns. For the utility metric, we train logistic regression [^7] https://scikit-learn.org/1.5/modules/generated/sklearn.linear_model.LogisticRegression.html, and XGB model [^8] https://xgboost.readthedocs.io/, reporting the average F1-score (F1), area under the receiver operating characteristic curve (AUC), and the accuracy score (ACC) for each dataset. For the histogram metric (HIST), we compute averages using bins of 20 and 50. The individual results for each metrics is shown in Appendix E. Additionally, for language model-based methods, we employ the perplexity metric to evaluate the model’s ability to generate accurate next tokens, focusing specifically on value perplexity (Value Perp) by masking out the perplexity of the value tokens.
Each model is run five times and four synthetic datasets generated per run. The size of the synthethic data is the same as the training data.
@@ 5.2 Non-DP Benchmarks
Table 2 highlights the effectiveness of LLMs using GPT-2 in comparison to other benchmark methods. GPT-2, fine-tuned over 10 epochs outperforms other models in terms of utility on the Adult dataset and achieves competitive results in terms of the histogram intersection metric. For the Airline dataset, the performance is also competitive compared to models such as CTGAN, TVAE, and VAE. These results clearly demonstrate the strong performance and capability of LLMs in this context. The Real data baseline represents the optimal performance achievable, calculated by evaluating the metrics using the real training data and comparing them against the real test data.
Adult Airline
Method F1 AUC ACC HIST F1 AUC ACC HIST
Real data 69.9$\pm$2 91.7$\pm$1 84.0$\pm$4 99.1$\pm$0 90.6$\pm$8 96.2$\pm$5 91.8$\pm$7 99.4$\pm$0
$\varepsilon=\infty$ CTGAN 59.5$\pm$6 88.5$\pm$0 80.2$\pm$3 91.2$\pm$1 87.2$\pm$3 94.7$\pm$2 88.9$\pm$3 94.4$\pm$1
TVAE 63.2$\pm$2 87.5$\pm$1 77.7$\pm$4 91.5$\pm$1 85.8$\pm$5 93.0$\pm$6 87.2$\pm$6 90.3$\pm$2
VAE 53.8$\pm$10 86.6$\pm$1 80.5$\pm$1 73.3$\pm$3 79.8$\pm$1 91.1$\pm$1 80.0$\pm$1 76.8$\pm$0
GPT-2 68.9$\pm$0 90.7$\pm$0 83.7$\pm$2 90.7$\pm$1 89.6$\pm$5 95.9$\pm$3 91.4$\pm$4 90.8$\pm$1
Table 2: Non-DP Benchmark. The Real data baseline represents the optimal achievable performance, determined by evaluating metrics using real training data compared to real test data. Utility metrics are reported as averages, with each model run five times and four synthetic datasets generated per run with standard deviation reported after $\pm$. The best value per column for each $\varepsilon$ is shown in bold while the second best value is underlined.

### v2

@@ 1 Introduction
In this work, we demonstrate that directly fine-tuning LLMs under DP constraints leads to sub-optimal performance. Our findings, shown in Figure 3, reveal that LLMs struggle to generate coherent and structured text when fine-tuned directly with DP. This challenge stems from inefficient privacy budget allocation to potentially non-private elements, such as the table column name “<key>” or non-functional tokens used to form a valid sentence for the LLM (e.g., “is ,”).
@@
We evaluated DP-2Stage on three datasets and observed improvements in utility metric (F1-score) by 12-25%, and marginal distribution metric (Hist) by 1-3% compared to direct DP fine-tuning (DP-Standard) demonstrating better utility and fidelity. Notably, DP-2Stage-U achieved up to 21× faster inference compared to DP-2Stage-O and DP-Standard. This speedup stems from its higher similarity to private data—both in key structure and value distribution—allowing the second-stage DP fine-tuning to focus less on learning structural patterns. Additionally, we observe that fine-tuning with DP while avoiding column shuffling, as used in  Borisov et al. (2023), yielded better performance under DP constraints but underperformed in Non-DP scenarios.
@@
  • Achieved 3–7% relative reductions in perplexity, 12–25% improvements in the utility metric (F1-score), and 1–3% relative improvements in the marginal distribution metric (Hist) compared to direct DP fine-tuning (DP-Standard). Each experiment was repeated five times per model, with four synthetic datasets generated per model to ensure robust evaluation.
@@ 2 Related Work
Language models have been extensively studied over the years, evolving from statistical models (Jelinek, 1998) and recurrent neural networks (Hochreiter & Schmidhuber, 1997) to the latest transformer-based architectures (Vaswani et al., 2017). These advancements, supported by the attention mechanisms and rich text datasets, have led to the emergence of large-scale language models (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023; Achiam et al., 2023), a new generation of language models. These models, together with various fine-tuning strategies (Hu et al., 2021; Zhou et al., 2022), have enabled several interesting applications (Borisov et al., 2023; Wang & Fritz, 2024). Borisov et al. (2023) proposed using pre-trained LLMs for synthesizing tabular data in a non-private setting by converting tabular data into text-like formats, significantly improving the performance and paving a new path toward LLM-based tabular generators. In light of the potential, our work apply these advanced models to a relatively unexplored domain: the generation of tabular data under differential privacy constraints. While previous studies (Yu et al., 2021; Li et al., 2022) have explored private fine-tuning of LLM, our focus specifically targets tabular data generation, a unique challenge that diverges from the broader scope of text comprehension. Recently, a concurrent work (Tran & Xiong, 2024) shared a similar spirit of employing two-stage training technique to DP fine-tune LLM for tabular data generation. However, this work does not provide insight into the impact of pseudo data used in the first stage and examines a different family of LLM and fine-tuning strategy compared to this work. Our focus is on GPT-2, as utilized in the pioneering work by Borisov et al. (2023), to gain deeper insights into how effective design choices in non-differentially private (Non-DP) settings translate to DP contexts. We offer a detailed analysis of pseudo-data choice, the impact of column shuffling on generation of coherent synthetic tabular data, identify the limitations of the two-stage approach, and discuss open challenges.
@@ 3.1 Language Models for Tabular Data
Let $f$ represent template, which defines the general pattern for organizing tabular data. Using GReaT serialization (Borisov et al., 2023), the template is expressed as $f=``\texttt{<key> is <value>,}^{\prime\prime}$, specifying how keys ($k\in{\mathcal{K}}$), values ($v\in{\mathcal{V}}$), and non-functional tokens ($c\in{\mathcal{C}}$, such as “is” and “,”) combine to form a record. An input dataset denoted as ${\mathcal{S}}$, is a realized instance of this $f$, instantiated with tokens $(k,v)$ from a record ${\bm{w}}=({\mathcal{K}},{\mathcal{V}})$. For example, with $k=\{\texttt{age}\}$ and $v=\{32\}$, ${\mathcal{S}}$ is instantiated as $f(k,v)=``\texttt{age is 32,}^{\prime\prime}$.
@@
Column shuffling involves randomly altering the order of column-aligned values within each batch during training. For example, given columns age, sex, and income with values 32, female, and >50k. A detailed illustration is provided in Figure 2. This mechanism happens at every iteration and has been shown to effectively prevent the model from relying on spurious dependency in Non-DP settings (Borisov et al., 2023). However, it complicates DP training due to gradient perturbations, often resulting in higher perplexity and different behavior compared to Non-DP models, as shown in Figure 3. Disabling shuffling fixes the order across iterations simplifying DP training.
@@
Figure 2: Illustration of column shuffling. The order of entries is permuted in each iteration. This mechanism happens at every iteration and has been shown to effectively prevent the model from relying on spurious dependency in Non-DP settings (Borisov et al., 2023). However, we find that it complicates DP training due to gradient perturbations, often resulting in higher perplexity compared to Non-DP models, as shown in Figure 3.
@@ 3.2 Differential Privacy
Figure 3: DP-2Stage (Ours) vs. Standard DP fine-tuning on the Adult dataset with $\varepsilon=1,\delta=10^{-5}$. DP-2Stage-O refers to the stage 2 model fine-tuned using out-distribution pseudo data (Airline dataset) in stage 1, while DP-2Stage-U is the stage 2 model fine-tuned using data sampled independently from a Uniform distribution, with statistics derived from the Adult dataset as the pseudo-dataset in stage 1. Perplexity results are displayed from left to right for all tokens, values, keys, and non-functional tokens (c.f. section 3.1). The top plot shows model trained without column shuffling, and the bottom shows model trained with column shuffling. Total and Value Perplexity for top and bottom plots have fixed y-axis for ease of comparison. The perplexity values are higher with column shuffling (bottom) than without (top).
@@ 4 DP-2Stage
While large language models (LLMs) demonstrate impressive capabilities in generating synthetic tabular data in non-differentially private (Non-DP) settings, as we will later show, fine-tuning them under DP constraints remains challenging for tabular data generation. As illustrated in Figure 3 (bottom), standard DP training, where the LLM model is directly fine-tuned using DPSGD (Abadi et al., 2016) (see Appendix A.1 for more details), often results in suboptimal performance in terms of the perplexity metric. We hypothesize that this is due to the inefficient allocation of the privacy budget during the learning of data structures. Models employing this approach struggle to reduce the perplexity of keys (e.g., column names) and other non-functional tokens, despite these elements typically posing minimal privacy risks. Consequently, applying DP to such components not only wastes the privacy budget but also diminishes the overall utility of the model.
@@ 5.1 Implementation Details
We utilized the GPT-2 (Radford et al., 2019) model from HuggingFace [^1] https://huggingface.co/ and integrated Opacus [^2] https://opacus.ai/ for differential privacy (DP) training, leveraging the BatchMemoryManager for efficient micro-batching. We fully fine-tuned the model with DP-Adam optimizer using a learning rate of $5e{-5}$ and a linear scheduler, setting the maximum gradient norm $C=1$ and a target delta $\delta=10^{-5}$. In the second stage of DP-2Stage, the parameter $\lambda$ was consistently set to 0.65 across all experiments. To convert the table-to-text, we used the GReaT serialization method (Borisov et al., 2023). Unlike GReaT, we maintained a fixed column ordering for each sampled dataset for the DP. We structure the dataset so that each row begins with the target column for that dataset. The Non-DP benchmark using GPT-2, the DP-Standard model, and the stage 2 models of DP-2Stage are all trained for 10 epochs, while the stage 1 model of DP-2Stage is trained for 5 epochs.
@@
To sample from the model, we condition on the target column for each dataset. For example, in the Adult dataset, where “income” is the target column, sampling begins with the prompt “income is ”, and generation continues until all columns have been sampled. We use a rejection sampling approach to discard incomplete generations. However, for the ablation results in Table 4 where column shuffling is enabled, rejection sampling proved to be very slow and sometimes failed to complete any generation. In these cases, we employed an imputation-based approach, as detailed in Appendix B.1.
@@
We compare our approach to Non-LLM-based methods under DP and Non-DP settings. For Non-DP baselines, we considered CTGAN, TVAE (Xu et al., 2019) and VAE (adaptation of TVAE using standardized preprocessing technique for the numerical columns i.e removing the mean and scaling to unit variance) while DP baselines include DP-CTGAN, DP-GAN, and DP-VAE. For CTGAN and TVAE, we used the implementation from the Synthetic Data Vault project [^3] https://github.com/sdv-dev/CTGAN, while DP-CTGAN and DP-GAN were implemented using the SmartNoise SDK [^4] https://github.com/opendp/smartnoise-sdk. The numerical column preprocessing was conducted with $\varepsilon=0.1$, while the remaining privacy budget was reserved for private training. For DP-VAE, we adapted the VAE model using SmartNoise preprocessing with $\varepsilon=0.1$.
All experiments were conducted using NVIDIA A100-SXM4-40GB GPUs. We train the models five times and generated four synthetic datasets per run. The size of the synthethic data is the same as the training data.
5.1.1 Datasets
We evaluated three tabular datasets: the Adult Income dataset, which contains over 30,000 samples and 15 attributes used to predict whether an individual’s annual income exceeds \$50,000, hosted on the UCI Machine Learning Repository [^5] https://archive.ics.uci.edu/dataset/2/adult; the Airline Passenger Satisfaction dataset, which includes over 100,000 samples and 24 attributes related to passenger satisfaction, available on Kaggle [^6] https://www.kaggle.com/datasets/teejmahal20/airline-passenger-satisfaction; and the Texas Hospital Discharge dataset [^7] https://www.dshs.texas.gov/thcic/, a large publicly available dataset provided by the Texas Department of State Health Services. We used the 2013 patient dataset that is preprocessed by (Stadler et al., 2022) and available on Github [^8] https://github.com/spring-epfl/synthetic_data_release/blob/master/data/texas.csv. Following  Afonja et al. (2023), we framed the task as a binary classification problem, predicting minor vs. major mortality risk. Unlike the more commonly benchmarked Adult dataset, the Airline and Texas dataset offers a fresh perspective on the ability of deep learning models to capture complex relationships within diverse tabular data and large sample sizes. Table 1 summarizes the statistics of the tabular data. Class Ratio represents the proportion of target class categories within each dataset. We provide more details in Appendix D.1.
# Train # Test # N # C # Total Class Ratio
Adult 30932 16858 6 9 15 76:24
Airline 103904 24976 19 5 24 57:43
Texas 60127 13978 7 11 18 81:19
@@ 5.1.1 Datasets
When training DP-2Stage-O on the Adult dataset, which is treated as the private dataset to be protected, we use the publicly available Airline and Texas datasets as out-of-distribution pseudo data for training the Stage 1 model. Similarly, for the Airline dataset, we use Adult and Texas, and for Texas, we use Adult and Airline as pseudo data.
5.1.2 Evaluation Metrics
We assess the quality of synthetic tabular datasets using a held-out test set and four key metrics inspired by prior work  (Xu et al., 2019; Afonja et al., 2023; Tao et al., 2021). These metrics are categorized into two groups: Utility and Fidelity.
Utility metrics evaluate how well synthetic data preserves patterns relevant to downstream machine learning tasks. Specifically, we use (i) Machine Learning Efficacy (ML Efficacy), which compares the predictive performance of models trained on synthetic data to those trained on real data.
Fidelity metrics assess the statistical similarity between real and synthetic data. (ii) The Normalized Histogram Intersection (HIST) measures how closely the marginal distributions of synthetic data align with those of real data. (iii) Pairwise Correlation Similarity Accuracy (CorAcc) evaluates how well synthetic data preserves pairwise column correlations from the real dataset. (iv) Pairwise Attribute Distribution Similarity (Pair) computes the similarity of all two-way attribute distributions by averaging histogram intersections, with numerical attributes discretized into bins.
By employing these metrics, we ensure that synthetic datasets not only support effective machine learning tasks but also maintain key statistical properties of the private data.
For ML Efficacy metric, we train logistic regression [^9] https://scikit-learn.org/1.5/modules/generated/sklearn.linear_model.LogisticRegression.html, and XGB model [^10] https://xgboost.readthedocs.io/, reporting the F1-score (F1), area under the receiver operating characteristic curve (AUC), and the accuracy score (ACC). For HIST and Pair, we compute averages using bins of 20 and 50. CorAcc is computed using Cramer’s V (with bias correction) for categorical columns, Correlation Ratio for numerical-categorical columns and Pearson Correlation Coefficient (absolute values) for the numerical columns.
Additionally, for language model-based methods, we evaluate the perplexity, specifically focusing on value perplexity (Value Perp), which measures the model’s ability to generate accurate next token while masking out the perplexity of the value tokens. More details about the metrics can be found in Appendix C.
@@ 5.2 Non-DP Benchmarks
Table 2 highlights the effectiveness of LLMs using GPT-2 in comparison to other benchmark methods. GPT-2, fine-tuned over 10 epochs outperforms other models in terms of utility on the Adult dataset and achieves competitive results for Pair and Hist metric. For the Airline and Texas dataset, the LLM either outperforms the GAN and VAE-based methods or achieves the second best. These results clearly demonstrate the strong performance and capability of LLMs in this context. The Real data baseline represents the optimal performance achievable, calculated by evaluating the metrics using the real training data and comparing them against the real test data.
Dataset Method F1 AUC ACC CorAcc Pair HIST
Adult Real data 69.9$\pm$2 91.7$\pm$1 84.0$\pm$4 97.3 97.5$\pm$0 99.1$\pm$0
$\varepsilon=\infty$ CTGAN 59.5$\pm$6 88.5$\pm$0 80.2$\pm$3 74.9$\pm$4 85.0$\pm$1 91.2$\pm$1
TVAE 63.2$\pm$2 87.5$\pm$1 77.7$\pm$4 76.6$\pm$2 84.5$\pm$1 91.5$\pm$1
VAE 53.8$\pm$10 86.6$\pm$1 80.5$\pm$1 65.3$\pm$2 60.2$\pm$2 73.3$\pm$3
GPT-2 68.9$\pm$0 90.7$\pm$0 83.7$\pm$2 79.9$\pm$1 83.7$\pm$0 90.7$\pm$1
Airline Real data 90.6$\pm$8 96.2$\pm$5 91.8$\pm$7 97.2 98.4$\pm$0 99.4$\pm$0
$\varepsilon=\infty$ CTGAN 87.2$\pm$3 94.7$\pm$2 88.9$\pm$3 84.4$\pm$1 89.3$\pm$1 94.4$\pm$1
TVAE 85.8$\pm$5 93.0$\pm$6 87.2$\pm$6 78.0$\pm$3 82.8$\pm$3 90.3$\pm$2
VAE 79.8$\pm$1 91.1$\pm$1 80.0$\pm$1 67.8$\pm$5 60.4$\pm$1 76.8$\pm$0
GPT-2 89.6$\pm$5 95.9$\pm$3 91.4$\pm$4 86.5$\pm$4 85.5$\pm$2 90.8$\pm$1
Texas Real data 86.6$\pm$2 98.8$\pm$0 95.0$\pm$1 96.3$\pm$0 98.9$\pm$0 99.5$\pm$0
$\varepsilon=\infty$ CTGAN 80.8$\pm$4 96.8$\pm$2 92.8$\pm$2 72.8$\pm$4 88.3$\pm$1 93.1$\pm$0
TVAE 76.7$\pm$7 95.3$\pm$3 90.9$\pm$3 61.2$\pm$4 63.4$\pm$7 93.5$\pm$1
VAE 74.3$\pm$3 95.6$\pm$1 90.6$\pm$2 51.5$\pm$7 57.5$\pm$1 89.3$\pm$1
GPT-2 83.9$\pm$2 98.6$\pm$0 93.6$\pm$1 72.8$\pm$1 82.8$\pm$3 94.9$\pm$0
Table 2: Non-DP Benchmark. The Real data baseline represents the optimal achievable performance, determined by evaluating metrics using real training data compared to real test data. F1, AUC and ACC are reported as averages of two ML Models (XGBoost and Logistic Regression). Pair, and Hist are reported as averages of two bin sizes (Bins 20 and 50). Each model run five times and four synthetic datasets generated per run with standard deviation reported after $\pm$. The best value per column for each $\varepsilon$ is shown in bold while the second best value is underlined.